OpenStudy (anonymous):

find inverse of y=x^2-1??

OpenStudy (tkhunny):

Swap and solve! What's holding you back?

OpenStudy (anonymous):

\[y \pm \sqrt{x-1}\]

OpenStudy (anonymous):

is that right?

OpenStudy (acacia):

y(x)=\[\pm \sqrt{y+1}\]

OpenStudy (anonymous):

thats not in my options

OpenStudy (acacia):

idk if mine is right /:

OpenStudy (anonymous):

theres one like that but instead its an x+1

OpenStudy (anonymous):

yes, should include \(x+1\)

OpenStudy (anonymous):

ah ok Thanx!!

OpenStudy (anonymous):

\[y=x^2-1\] \[x=y^2-1\] \[x+1=y^2\] \[\pm\sqrt{x+1}=y\]

OpenStudy (anonymous):

Substitute it in to ceck i.e ... sub in \[(\sqrt{x+1)}^2-1\] and simplify

OpenStudy (anonymous):

yeah thats the one i have in my options